# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 2.1. Riemannian homogeneous spaces
Let $M=\mathsf{G}/\mathsf{K}$ be a Riemannian homogeneous space and $o=e\mathsf{K}$. We denote by $\mathsf{G}_{p}$ the isotropy subgroup of $p$, that is, the set of all elements of $\mathsf{G}$ that fix $p$. For our purposes we will assume that $\mathsf{G}$ is connected and the action of $\mathsf{G}$ on $M$ is almost effective, in the sense that $\bigcap_{p\in M}\mathsf{G}_{p}$ is a discrete subgroup of $\mathsf{G}$. The (Lie) group of isometries of $M$ will be written as $I(M)$, and $I^{0}(M)\subseteq I(M)$ denotes its identity component. Lie groups will be denoted by uppercase letters and their Lie algebras will be denoted by their corresponding lowercase gothic letters.
@@
On the other hand, we have the Levi-Civita connection $\nabla$. We can also consider the tensor $D:=\nabla-\nabla^{c}$. It turns out that $D$ is a $\mathsf{G}$-invariant tensor defined on $M=\mathsf{G}/\mathsf{K}$ called the difference tensor. Notice that $U=0$ holds if and only if the operators $A_{X}\colon Y\in\mathfrak{p}\to[X,Y]_{\mathfrak{p}}$ are skew-symmetric for all $X\in\mathfrak{p}$. In particular, if $D$ is an element of $\operatorname{Hom}_{\mathsf{K}}\left(\wedge^{2}\mathfrak{p},\mathfrak{p}\right)$, then the reductive decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ is naturally reductive.
Following [39, Chapter X], we can express $\nabla$ and $D$ at the base point $o\in M$ in terms of the Lie bracket of $\mathfrak{g}$ and the metric of $M$ in the following way:
@@
where $X,Y,Z\in\mathfrak{p}$. We will also need to compute the covariant derivative $\nabla R$ of the curvature tensor. As it is shown in [48, Remark 2.5], $\nabla R$ is given at $o$ by
@@
An important class of Riemannian submersions involving homogeneous spaces is that of homogeneous fibrations, which we briefly describe. The main reference is [26]. Let $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$ be a chain of inclusions of compact connected subgroups of the connected Lie group $\mathsf{G}$, and endow $\mathsf{G}/\mathsf{K}$ with a $\mathsf{G}$-invariant Riemannian metric. Then there exists a left-invariant metric on $\mathsf{G}$ that is also right $\mathsf{K}$-invariant, and a $\mathsf{G}$-invariant metric on $\mathsf{G}/\mathsf{H}$ that makes the canonical projection $\pi\colon\mathsf{G}/\mathsf{H}\to\mathsf{G}/\mathsf{K}$ a Riemannian submersion with totally geodesic fibers isometric to $\mathsf{K}/\mathsf{H}$. We denote by $\mathcal{V}$ and $\mathcal{H}$ the vertical and horizontal distributions associated with the Riemannian submersion $\pi$. We will be concerned with the case that $\mathsf{G}$ is a compact group with a bi-invariant metric and the homogeneous spaces $F=\mathsf{K}/\mathsf{H}$, $M=\mathsf{G}/\mathsf{H}$ and $B=\mathsf{G}/\mathsf{K}$ are endowed with the corresponding normal homogeneous metrics. If $V$ is a Euclidean vector space and $W\subseteq V$ is a vector subspace, we denote by $V\ominus W$ the orthogonal complement of $W$ in $V$. In this case, the tangent space $T_{o}M$ at $o=e\mathsf{H}$ is identified with $\mathfrak{p}=\mathfrak{g}\ominus\mathfrak{h}$, so the vertical and horizontal subspaces at $o$ are $\mathcal{V}_{o}=\mathfrak{k}\ominus\mathfrak{h}$ and $\mathcal{H}_{o}=\mathfrak{p}\ominus\mathfrak{k}$. It turns out that the distributions $\mathcal{V}$ and $\mathcal{H}$ are $\mathsf{G}$-invariant in the sense that for every $p\in M$ and $g\in\mathsf{G}$, we have $g_{*p}\mathcal{V}_{p}=\mathcal{V}_{g\cdot p}$ and $g_{*p}\mathcal{H}_{p}=\mathcal{H}_{g\cdot p}$.
A submanifold $N\subseteq M$ is said to be extrinsically homogeneous with respect to the presentation $M=\mathsf{G}/\mathsf{K}$ if there exists a Lie subgroup $\mathsf{S}\subseteq\mathsf{G}$ such that $N$ is an orbit of $\mathsf{S}$. More generally, a submanifold $N$ of a Riemannian manifold $M$ is extrinsically homogeneous if it is an orbit of a Lie subgroup $\mathsf{S}\subseteq I(M)$. In the following lemma, we provide a formula for the second fundamental form for an extrinsically homogeneous submanifold of a homogeneous space. This formula was originally derived by Solonenko in the case of Riemannian symmetric spaces, see [53, Proposition 2.2.43]. See also [2, Proposition 2.2] for an alternative expression.
@@
Thus, projecting to the normal space and using the fact that the second fundamental form is symmetric, we obtain that
@@ 2.2. Homogeneous nearly Kähler manifolds
In this section we present the ambient spaces that we will work with throughout the rest of this article. The classification of homogeneous nearly Kähler manifolds in dimension six was done by Butruille [12]. Indeed, every simply connected Riemannian manifold satisfying the previous conditions is homothetic to either the sphere $\mathsf{S}^{6}$, the complex projective space $\mathbb{C}\mathsf{P}^{3}$, the flag manifold $\mathsf{F}(\mathbb{C}^{3})$, and $\mathsf{S}^{3}\times\mathsf{S}^{3}$. Since $\mathsf{S}^{6}$ carries its natural round metric, the totally geodesic submanifolds of $\mathsf{S}^{6}$ are well-known, so we will only focus on the other three spaces. It turns out that these manifolds are examples of $3$-symmetric spaces. The main reference for the description of $3$-symmetric spaces is [25].
@@
Any 3-symmetric space is automatically homogeneous [22, Theorem 4.8]. Conversely, one can construct a $3$-symmetric space in terms of algebraic data. Indeed, let $\mathsf{G}$ be a connected Lie group and $\mathsf{K}$ a closed subgroup of $\mathsf{G}$. Assume that there exists an automorphism $\Theta\colon\mathsf{G}\to\mathsf{G}$ of order three such that $\mathsf{G}^{\Theta}_{0}\subseteq\mathsf{K}\subseteq\mathsf{G}^{\Theta}$, where $\mathsf{G}^{\Theta}$ is the fixed point set of $\Theta$ and $\mathsf{G}^{\Theta}_{0}$ is its identity component. It turns out that $M=\mathsf{G}/\mathsf{K}$ is a reductive homogeneous space in a way that $\theta=\Theta_{*}$ preserves the reductive complement. Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be a reductive decomposition of $\mathfrak{g}$ satisfying $\theta\mathfrak{p}=\mathfrak{p}$. Then, any inner product on $\mathfrak{p}$ that is invariant under $\operatorname{Ad}(\mathsf{K})$ and $\theta$ gives rise to a $\mathsf{G}$-invariant metric on $M$ that turns $M$ into a $3$-symmetric space, where the isometry of order three at $o=e\mathsf{K}$ is given by $s_{o}(x\mathsf{K})=\Theta(x)\mathsf{K}$. We say that $(\mathsf{G},\mathsf{K},\Theta)$ is the triple associated with the $3$-symmetric space $M$. The corresponding almost complex structure is the $\mathsf{G}$-invariant tensor field $J$ defined at $o$ by (3). By [22, Proposition 5.6], the almost Hermitian manifold $(M,J)$ is nearly Kähler if and only if the decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ is naturally reductive.
@@ 2.2.1. The complex projective space $\mathbb{C}\mathsf{P}^{3}$
Let $E_{ij}\in\mathfrak{gl}(2,\mathbb{H})$ be the elementary matrix which has all components equal to zero except for the $(i,j)$ component, which is one. We will use the orthonormal basis $\{e_{1},\dots,e_{6}\}$ of $\mathfrak{p}$ defined by
@@
The isotropy representation allows us to decompose $\mathfrak{p}$ as the direct sum of two irreducible submodules $\mathfrak{p}_{1}=\operatorname{span}\{e_{1},e_{2}\}$ and $\mathfrak{p}_{2}=\operatorname{span}\{e_{3},\dots,e_{6}\}$. Indeed, the subrepresentation $\mathfrak{p}_{1}$ of $\mathsf{U}(1)\times\mathsf{Sp}(1)$ is isomorphic to the representation $\mathbb{C}$ with the action given by $(\lambda,\mu)\cdot z=\lambda^{2}z$, whereas $\mathfrak{p}_{2}$ is isomorphic to the representation $\mathbb{R}^{4}$ of $\mathsf{U}(1)\times\mathsf{Sp}(1)$ under $(\lambda,\mu)x=\mu x\bar{\lambda}$. In particular, both subrepresentations are transitive on the unit sphere of each $\mathfrak{p}_{i}$.
@@ 2.2.2. The flag manifold $\mathsf{F}(\mathbb{C}^{3})$
Recall that a full flag in $\mathbb{C}^{3}$ is a chain $0=V_{0}\subseteq V_{1}\subseteq V_{2}\subseteq V_{3}=\mathbb{C}^{3}$ (also denoted by $(V_{1},V_{2})$) of subspaces such that $\dim_{\mathbb{C}}V_{k}=k$ for each $k$. We denote by $\mathsf{F}(\mathbb{C}^{3})$ the space of all flags in $\mathbb{C}^{3}$, which is naturally identified with the quotient of the Stiefel manifold of orthonormal bases of $\mathbb{C}^{3}$ under the standard action of $\mathsf{U}(1)^{3}$. The group $\mathsf{G}=\mathsf{SU}(3)$ acts transitively on $\mathsf{F}(\mathbb{C}^{3})$, and if $o$ is the standard flag $0\subseteq\mathbb{C}e_{1}\subseteq\operatorname{span}\{e_{1},e_{2}\}\subseteq\mathbb{C}^{3}$, its isotropy subgroup is the maximal torus $\mathsf{T}^{2}$ of diagonal matrices in $\mathsf{SU}(3)$, so we have $\mathsf{F}(\mathbb{C}^{3})=\mathsf{SU}(3)/\mathsf{T}^{2}$.
@@
Once again, the $E_{ij}$ denote elementary $3\times 3$ matrices. It is easy to check that the tangent space splits as the direct sum of irreducible submodules $\mathfrak{p}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}\oplus\mathfrak{p}_{3}$, where each $\mathfrak{p}_{k}=\operatorname{span}\{e_{2k-1},e_{2k}\}$ is isomorphic to $\mathbb{C}$. To be more precise, if $g=\operatorname{diag}(e^{ix},e^{iy},e^{-i(x+y)})$ is an arbitrary element of $\mathsf{T}^{2}$, then $\operatorname{Ad}(g)$ acts on $\mathfrak{p_{1}}$ as multiplication by $e^{i(x-y)}$, on $\mathfrak{p}_{2}$ as multiplication by $e^{i(x+2y)}$, and on $\mathfrak{p}_{3}$ as multiplication by $e^{i(2x+y)}$. Note that $\mathfrak{p}_{1}$, $\mathfrak{p}_{2}$ and $\mathfrak{p}_{3}$ are pairwise nonisomorphic as representations of $\mathsf{T}^{2}$. Furthermore, if $g\in\mathsf{U}(3)$ is a permutation matrix, then the map $a\mathsf{T}^{2}\to gag^{-1}\mathsf{T}^{2}$ is an isometry fixing $o$ and whose differential at $o$ permutes the irreducible submodules of $\mathfrak{p}$, and every permutation of these submodules can be achieved in this way. For example, the transposition $(1,2)$ interchanges $\mathfrak{p}_{2}$ and $\mathfrak{p}_{3}$ and the cycle $(1,2,3)$ acts as the cycle $(\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3})$.
@@
We now determine the full isometry group of $\mathsf{F}(\mathbb{C}^{3})$. This computation was done by Shankar in [52] when $\mathsf{F}(\mathbb{C}^{3})$ carries a metric of positive sectional curvature. However, the homogeneous metric that we are considering in $\mathsf{F}(\mathbb{C}^{3})$ does not have positive sectional curvature. In our case, we may calculate the isometry group of $\mathsf{F}(\mathbb{C}^{3})$ via the following approach (based on the proof of [49, §4, Proposition 6 and §16, Theorem 3]). Firstly, the effectivized version of the presentation $\mathsf{SU}(3)/\mathsf{T}^{2}$ is $\mathsf{PSU(3)}/(\mathsf{T}^{2}/\mathbb{Z}_{3})$, and we may apply [59, Theorem 5.1] to conclude that $I^{0}(\mathsf{F}(\mathbb{C}^{3}))=\mathsf{PSU}(3)$. As for the group of components $I(\mathsf{F}(\mathbb{C}^{3}))/I^{0}(\mathsf{F}(\mathbb{C}^{3}))$, since the flag manifold is simply connected, this group is equal to $\mathsf{H}/\mathsf{H}^{0}$, where $\mathsf{H}$ is the isotropy subgroup of $I(\mathsf{F}(\mathbb{C}^{3}))$ at $o$ and $\mathsf{H}^{0}$ its identity component. This follows from the long exact sequence of homotopy groups associated with the fibration $\mathsf{H}\hookrightarrow I(\mathsf{F}(\mathbb{C}^{3}))\to\mathsf{F}(\mathbb{C}^{3})$. Now, the conjugation map $C\colon\mathsf{H}\to\mathsf{Aut}(\mathsf{PSU}(3),\mathsf{T}^{2}/\mathbb{Z}_{3})=\{\varphi\in\mathsf{Aut}(\mathsf{PSU}(3)):\varphi\text{ preserves }\mathsf{T}^{2}/\mathbb{Z}_{3}\}$ is injective by [52, Proposition 1.7]. In addition, any $\varphi\in\mathsf{Aut}(\mathsf{PSU}(3),\mathsf{T}^{2}/\mathbb{Z}_{3})$ descends to a diffeomorphism $\bar{\varphi}$ of $\mathsf{F}(\mathbb{C}^{3})$, which is actually an isometry because $\varphi$ preserves the Killing form and the metric on the flag manifold is induced by it. It is easy to show that $\bar{\varphi}\in\mathsf{H}$ and $C(\bar{\varphi})=\varphi$, so $C$ is an isomorphism and $\mathsf{H}=\mathsf{Aut}(\mathsf{PSU}(3),\mathsf{T}^{2}/\mathbb{Z}_{3})$. Now, since $\mathsf{Aut}(\mathsf{PSU}(3))=\operatorname{Ad}(\mathsf{PSU}(3))\rtimes\mathbb{Z}_{2}$, where the outer $\mathbb{Z}_{2}$ is generated by complex conjugation, the computation of $\mathsf{H}/\mathsf{H}^{0}$ reduces to that of $\frac{\mathsf{N}_{\mathsf{PSU}(3)}(\mathsf{T}^{2}/\mathbb{Z}_{3})}{\mathsf{Z}_{\mathsf{PSU}(3)}(\mathsf{T}^{2}/\mathbb{Z}_{3})}\rtimes\mathbb{Z}_{2}$. The first factor is merely the Weyl group $\mathsf{W}(\mathsf{PSU}(3))=\mathfrak{S}_{3}$, so we have obtained $\mathsf{H}/\mathsf{H}^{0}=\mathfrak{S}_{3}\rtimes\mathbb{Z}_{2}$, and we conclude that the full isometry group is $I(\mathsf{F}(\mathbb{C}^{3}))=\mathsf{PSU}(3)\rtimes(\mathfrak{S}_{3}\rtimes\mathbb{Z}_{2})$.

### v3

@@ 2.1. Riemannian homogeneous spaces
Let $M=\mathsf{G}/\mathsf{K}$ be a Riemannian homogeneous space and $o=e\mathsf{K}$. We denote by $\mathsf{G}_{p}$ the isotropy subgroup of $p$, that is, the set of all elements of $\mathsf{G}$ that fix $p$. For our purposes we assume that $\mathsf{G}$ is connected and the action of $\mathsf{G}$ on $M$ is almost effective, in the sense that $\bigcap_{p\in M}\mathsf{G}_{p}$ is a discrete subgroup of $\mathsf{G}$. The (Lie) group of isometries of $M$ is written as $I(M)$, and $I^{0}(M)\subseteq I(M)$ denotes its identity component. Throughout this article, Lie groups are denoted by uppercase letters and their Lie algebras are denoted by their corresponding lowercase gothic letters.
@@
On the other hand, we have the Levi-Civita connection $\nabla$. We can also consider the tensor $D:=\nabla-\nabla^{c}$. It turns out that $D$ is a $\mathsf{G}$-invariant tensor defined on $M=\mathsf{G}/\mathsf{K}$ called the difference tensor. Notice that $U=0$ holds if and only if the operators $A_{X}\colon Y\in\mathfrak{p}\to[X,Y]_{\mathfrak{p}}$ are skew-symmetric for all $X\in\mathfrak{p}$. In particular, if $D$ is an element of $\operatorname{Hom}_{\mathsf{K}}\bigl{(}\bigwedge^{2}\mathfrak{p},\mathfrak{p}\bigr{)}$, then the reductive decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ is naturally reductive.
Following [38, Chapter X], we can express $\nabla$ and $D$ at the base point $o\in M$ in terms of the Lie bracket of $\mathfrak{g}$ and the metric of $M$ in the following way:
@@
where $X,Y,Z\in\mathfrak{p}$. We also need to compute the covariant derivative $\nabla R$ of the curvature tensor. As shown in [46, Remark 2.5], the tensor $\nabla R$ is given at $o$ by
@@
An important class of Riemannian submersions involving homogeneous spaces is that of homogeneous fibrations, which we briefly describe. The main reference is [25]. Let $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$ be a chain of inclusions of compact connected subgroups of the connected Lie group $\mathsf{G}$, and endow $\mathsf{G}/\mathsf{K}$ with a $\mathsf{G}$-invariant Riemannian metric. Then there exists a left-invariant metric on $\mathsf{G}$ that is also right $\mathsf{K}$-invariant, and a $\mathsf{G}$-invariant metric on $\mathsf{G}/\mathsf{H}$ that makes the canonical projection $\pi\colon\mathsf{G}/\mathsf{H}\to\mathsf{G}/\mathsf{K}$ a Riemannian submersion with totally geodesic fibers isometric to $\mathsf{K}/\mathsf{H}$. We denote by $\mathcal{V}$ and $\mathcal{H}$ the vertical and horizontal distributions associated with the Riemannian submersion $\pi$. We are concerned with the case that $\mathsf{G}$ is a compact group with a bi-invariant metric and the homogeneous spaces $F=\mathsf{K}/\mathsf{H}$, $M=\mathsf{G}/\mathsf{H}$ and $B=\mathsf{G}/\mathsf{K}$ are endowed with the corresponding normal homogeneous metrics. If $V$ is a Euclidean vector space and $W\subseteq V$ is a vector subspace, we denote by $V\ominus W$ the orthogonal complement of $W$ in $V$. In this case, the tangent space $T_{o}M$ at $o=e\mathsf{H}$ is identified with $\mathfrak{p}=\mathfrak{g}\ominus\mathfrak{h}$, so the vertical and horizontal subspaces at $o$ are $\mathcal{V}_{o}=\mathfrak{k}\ominus\mathfrak{h}$ and $\mathcal{H}_{o}=\mathfrak{p}\ominus\mathfrak{k}$. It turns out that the distributions $\mathcal{V}$ and $\mathcal{H}$ are $\mathsf{G}$-invariant in the sense that for every $p\in M$ and $g\in\mathsf{G}$, we have $g_{*p}\mathcal{V}_{p}=\mathcal{V}_{g\cdot p}$ and $g_{*p}\mathcal{H}_{p}=\mathcal{H}_{g\cdot p}$.
A submanifold $N\subseteq M$ is said to be extrinsically homogeneous with respect to the presentation $M=\mathsf{G}/\mathsf{K}$ if there exists a Lie subgroup $\mathsf{S}\subseteq\mathsf{G}$ such that $N$ is an orbit of $\mathsf{S}$. More generally, a submanifold $N$ of a Riemannian manifold $M$ is extrinsically homogeneous if it is an orbit of a Lie subgroup $\mathsf{S}\subseteq I(M)$. In the following lemma, we provide a formula for the second fundamental form for an extrinsically homogeneous submanifold of a homogeneous space. This formula was originally derived by Solonenko in the case of Riemannian symmetric spaces, see [52, Proposition 2.2.43]. See also [2, Proposition 2.2] for an alternative expression.
@@
Thus, projecting onto the normal space and using the fact that the second fundamental form is symmetric, we obtain that
@@ 2.2. Homogeneous nearly Kähler manifolds
In this section we present the ambient spaces that we work with throughout the rest of this article. The classification of homogeneous nearly Kähler manifolds in dimension six was done by Butruille [12]. Indeed, every simply connected Riemannian manifold satisfying the previous conditions is homothetic to either the sphere $\mathsf{S}^{6}$, the complex projective space $\mathbb{C}\mathsf{P}^{3}$, the flag manifold $\mathsf{F}(\mathbb{C}^{3})$, and $\mathsf{S}^{3}\times\mathsf{S}^{3}$. Since $\mathsf{S}^{6}$ carries its natural round metric, its totally geodesic submanifolds are well-known, so we only need to focus on the other three spaces. It turns out that these manifolds are examples of $3$-symmetric spaces. The main reference for the description of $3$-symmetric spaces is [59].
@@
Any connected 3-symmetric space is automatically homogeneous [22, Theorem 4.8]. Conversely, one can construct a $3$-symmetric space in terms of algebraic data. Indeed, let $\mathsf{G}$ be a connected Lie group and $\mathsf{K}$ a closed subgroup of $\mathsf{G}$. Assume that there exists an automorphism $\Theta\colon\mathsf{G}\to\mathsf{G}$ of order three such that $\mathsf{G}^{\Theta}_{0}\subseteq\mathsf{K}\subseteq\mathsf{G}^{\Theta}$, where $\mathsf{G}^{\Theta}$ is the fixed point set of $\Theta$ and $\mathsf{G}^{\Theta}_{0}$ is its identity component. It turns out that $M=\mathsf{G}/\mathsf{K}$ is a reductive homogeneous space in such a way that $\theta=\Theta_{*}$ preserves the reductive complement. Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be a reductive decomposition of $\mathfrak{g}$ satisfying $\theta\mathfrak{p}=\mathfrak{p}$. Then, any inner product on $\mathfrak{p}$ that is invariant under $\operatorname{Ad}(\mathsf{K})$ and $\theta$ gives rise to a $\mathsf{G}$-invariant metric on $M$ that turns $M$ into a $3$-symmetric space, where the isometry of order three at $o=e\mathsf{K}$ is given by $s_{o}(x\mathsf{K})=\Theta(x)\mathsf{K}$. We say that $(\mathsf{G},\mathsf{K},\Theta)$ is the triple associated with the $3$-symmetric space $M$. The corresponding almost complex structure is the $\mathsf{G}$-invariant tensor field $J$ defined at $o$ by (3). By [22, Proposition 5.6], the almost Hermitian manifold $(M,J)$ is nearly Kähler if and only if the decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ is naturally reductive.
@@ 2.2.1. The complex projective space $\mathbb{C}\mathsf{P}^{3}$
Let $E_{ij}\in\mathfrak{gl}(2,\mathbb{H})$ be the elementary matrix which has all components equal to zero except for the $(i,j)$ component, which is one. We use the orthonormal basis $\{e_{1},\dots,e_{6}\}$ of $\mathfrak{p}$ defined by
@@
The isotropy representation allows us to decompose $\mathfrak{p}$ as the direct sum of two irreducible submodules $\mathfrak{p}_{1}=\operatorname{span}\{e_{1},e_{2}\}$ and $\mathfrak{p}_{2}=\operatorname{span}\{e_{3},\dots,e_{6}\}$. Indeed, the subrepresentation $\mathfrak{p}_{1}$ of $\mathsf{U}(1)\times\mathsf{Sp}(1)$ is isomorphic to the representation $\mathbb{C}$ with the action given by $(\lambda,\mu)\cdot z=\lambda^{2}z$, whereas $\mathfrak{p}_{2}$ is isomorphic to the representation $\mathbb{R}^{4}$ of $\mathsf{U}(1)\times\mathsf{Sp}(1)$ under $(\lambda,\mu)\cdot x=\mu x\bar{\lambda}$. In particular, the group $\mathsf{U}(1)\times\mathsf{Sp}(1)$ acts transitively on the unit sphere of each $\mathfrak{p}_{i}$.
@@ 2.2.2. The flag manifold $\mathsf{F}(\mathbb{C}^{3})$
Recall that a full flag in $\mathbb{C}^{3}$ is a chain $0=V_{0}\subseteq V_{1}\subseteq V_{2}\subseteq V_{3}=\mathbb{C}^{3}$ (also denoted by $(V_{1},V_{2})$) of subspaces such that $\dim_{\mathbb{C}}V_{k}=k$ for each $k$. We denote by $\mathsf{F}(\mathbb{C}^{3})$ the space of all flags in $\mathbb{C}^{3}$, which is naturally identified with the quotient of the Stiefel manifold of orthonormal bases of $\mathbb{C}^{3}$ under the standard action of $\mathsf{U}(1)^{3}$. The group $\mathsf{G}=\mathsf{SU}(3)$ acts transitively on $\mathsf{F}(\mathbb{C}^{3})$, and if $o$ is the standard flag $0\subseteq\mathbb{C}e_{1}\subseteq\operatorname{span}\{e_{1},e_{2}\}\subseteq\mathbb{C}^{3}$, its isotropy subgroup is the maximal torus $\mathsf{K}=\mathsf{T}^{2}$ of diagonal matrices in $\mathsf{SU}(3)$, so we have $\mathsf{F}(\mathbb{C}^{3})=\mathsf{SU}(3)/\mathsf{T}^{2}$.
@@
Once again, the $E_{ij}$ denote elementary $3\times 3$ matrices. It is easy to check that the tangent space splits as the direct sum of irreducible submodules $\mathfrak{p}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}\oplus\mathfrak{p}_{3}$, where each $\mathfrak{p}_{k}=\operatorname{span}\{e_{2k-1},e_{2k}\}$ is isomorphic to $\mathbb{C}$. To be more precise, if $g=\operatorname{diag}\bigl{(}e^{ix},e^{iy},e^{-i(x+y)}\bigr{)}$ is an arbitrary element of $\mathsf{T}^{2}$, then $\operatorname{Ad}(g)$ acts on $\mathfrak{p_{1}}$ as multiplication by $e^{i(x-y)}$, on $\mathfrak{p}_{2}$ as multiplication by $e^{i(x+2y)}$, and on $\mathfrak{p}_{3}$ as multiplication by $e^{i(2x+y)}$. Note that $\mathfrak{p}_{1}$, $\mathfrak{p}_{2}$ and $\mathfrak{p}_{3}$ are pairwise nonisomorphic as representations of $\mathsf{T}^{2}$. Furthermore, if $g\in\mathsf{U}(3)$ is a permutation matrix, then the map $a\mathsf{T}^{2}\to gag^{-1}\mathsf{T}^{2}$ is an isometry fixing $o$ and whose differential at $o$ permutes the irreducible submodules of $\mathfrak{p}$, and every permutation of these submodules can be achieved in this way. For example, the transposition $(1,2)$ interchanges $\mathfrak{p}_{2}$ and $\mathfrak{p}_{3}$ and the cycle $(1,2,3)$ acts as the cycle $(\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3})$.
@@
We now determine the full isometry group of $\mathsf{F}(\mathbb{C}^{3})$. This computation was done by Shankar in [51] when $\mathsf{F}(\mathbb{C}^{3})$ carries a metric of positive sectional curvature. However, the homogeneous metric that we are considering in $\mathsf{F}(\mathbb{C}^{3})$ does not have positive sectional curvature. In our case, we may calculate the isometry group of $\mathsf{F}(\mathbb{C}^{3})$ via the following approach (based on the proof of [47, §4, Proposition 6 and §16, Theorem 3]). Firstly, the effectivized version of the presentation $\mathsf{SU}(3)/\mathsf{T}^{2}$ is $\mathsf{PSU(3)}/(\mathsf{T}^{2}/\mathbb{Z}_{3})$, and we may apply [58, Theorem 5.1] to conclude that $I^{0}(\mathsf{F}(\mathbb{C}^{3}))=\mathsf{PSU}(3)$. As for the group of components $I(\mathsf{F}(\mathbb{C}^{3}))/I^{0}(\mathsf{F}(\mathbb{C}^{3}))$, since the flag manifold is simply connected, this group is equal to $\mathsf{H}/\mathsf{H}^{0}$, where $\mathsf{H}$ is the isotropy subgroup of $I(\mathsf{F}(\mathbb{C}^{3}))$ at $o$ and $\mathsf{H}^{0}$ its identity component. This follows from the long exact sequence of homotopy groups associated with the fibration $\mathsf{H}\hookrightarrow I(\mathsf{F}(\mathbb{C}^{3}))\to\mathsf{F}(\mathbb{C}^{3})$. Now, the conjugation map $C\colon\mathsf{H}\to\mathsf{Aut}(\mathsf{PSU}(3),\mathsf{T}^{2}/\mathbb{Z}_{3})=\{\varphi\in\mathsf{Aut}(\mathsf{PSU}(3)):\varphi\text{ preserves }\mathsf{T}^{2}/\mathbb{Z}_{3}\}$ is injective by [51, Proposition 1.7]. In addition, any $\varphi\in\mathsf{Aut}(\mathsf{PSU}(3),\mathsf{T}^{2}/\mathbb{Z}_{3})$ descends to a diffeomorphism $\bar{\varphi}$ of $\mathsf{F}(\mathbb{C}^{3})$, which is actually an isometry because $\varphi$ preserves the Killing form and the metric on the flag manifold is induced by it. It is easy to show that $\bar{\varphi}\in\mathsf{H}$ and $C(\bar{\varphi})=\varphi$, so $C$ is an isomorphism and $\mathsf{H}=\mathsf{Aut}(\mathsf{PSU}(3),\mathsf{T}^{2}/\mathbb{Z}_{3})$. Now, since $\mathsf{Aut}(\mathsf{PSU}(3))=\operatorname{Ad}(\mathsf{PSU}(3))\rtimes\mathbb{Z}_{2}$, where the outer $\mathbb{Z}_{2}$ is generated by complex conjugation, the computation of $\mathsf{H}/\mathsf{H}^{0}$ reduces to that of $\frac{\mathsf{N}_{\mathsf{PSU}(3)}(\mathsf{T}^{2}/\mathbb{Z}_{3})}{\mathsf{Z}_{\mathsf{PSU}(3)}(\mathsf{T}^{2}/\mathbb{Z}_{3})}\rtimes\mathbb{Z}_{2}$. The first factor is merely the Weyl group $\mathsf{W}(\mathsf{PSU}(3))=\mathfrak{S}_{3}$, so we have obtained $\mathsf{H}/\mathsf{H}^{0}=\mathfrak{S}_{3}\rtimes\mathbb{Z}_{2}$, and we conclude that the full isometry group is $I(\mathsf{F}(\mathbb{C}^{3}))=\mathsf{PSU}(3)\rtimes(\mathfrak{S}_{3}\rtimes\mathbb{Z}_{2})$.